# Parameter estimation for Gibbs distributions

### v9

@@ 1 Introduction
A central problem in computational statistics is to convert a procedure for sampling combinatorial objects into a procedure for counting those objects, and vice versa. We will consider sampling algorithms for Gibbs distributions. Formally, given a real-valued function $H(\cdot)$ over a finite set $\Omega$, the Gibbs distribution is the family of distributions $\mu^{\Omega}_{\beta}$ over $\Omega$, parameterized by $\beta$ over an interval $[\beta_{\min},\beta_{\max}]$, of the form
@@
The normalizing constant $Z(\beta)=\sum_{\omega\in\Omega}e^{\beta H(\omega)}$ is called the partition function. These distributions appear in a number of sampling algorithms; they also frequently occur in physics, where the parameter $-\beta$ corresponds to the inverse temperature, and $H(\omega)$ is called the Hamiltonian of the system.
@@
There is an associated probability distribution we call the gross Gibbs distribution $\mu_{\beta}(x)$ over $\mathbb{R}$ given by
@@
Note that, for the gross Gibbs distributions, the counts $c_{x}$ can only be determined up to scaling. When $c_{0}>0$, the value $\beta=-\infty$ is allowed and meaningful; in this case, we have $Z(-\infty)=c_{0}$ and $\mu_{-\infty}$ is simply the distribution which is equal to 00 with probability one. In particular, for $\beta_{\min}=-\infty$ we have $Q(\beta)=Z(\beta)/c_{0}$.
@@
To avoid degenerate cases, we always assume $n,q\geq 2$ and $\varepsilon,\gamma\in(0,\tfrac{1}{2})$. If upper bounds $n$ and/or $q$ are not available directly, they can often be estimated by simple algorithms (up to constant factors), or can be guessed by exponential back-off strategies.
@@ 1.1 Our contributions
Our main contribution is a black-box algorithm to estimate counts. To explain this, we must address two technical issues. First, since the counts can only be recovered up to scaling in the black-box setting, some (arbitrary) normalization must be chosen. We use a convenient normalization $\pi(x)$ defined as:
@@
With these two provisos, we can define ${P^{\delta,\varepsilon}_{\tt count}}$ for parameters $\delta,\varepsilon\in(0,1)$ to be the problem of finding a vector $\hat{\pi}\in[0,1]^{\mathcal{F}}$ with the following two properties:
@@
In particular, if $\Delta(x)\geq\delta$, then ${P^{\delta,\varepsilon}_{\tt count}}$ provides a $(1\pm 3\varepsilon))$ relative approximation to $\pi(x)$. We emphasize here that when the algorithm succeeds (which should have probability at least $1-\gamma$), this guarantee holds simultaneously for all values $x$. We develop three main algorithmic results:
@@
We also show lower bounds for black-box algorithms to solve ${P^{\delta,\varepsilon}_{\tt count}}$ by sampling the gross Gibbs distribution. We summarize these results here as follows:
@@
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
@@
Some theoretical count-estimation algorithms have been considered for combinatorial problems depending on specific properties of the Gibbs distribution, see e.g. [18, 9]. In physical application, the Wang-Landau (WL) algorithm [26] is a popular heuristic to estimate counts; this uses a completely different methodology from our algorithm, based on a random walk on $\mathcal{F}$ with a running count estimate $\hat{c}$. As discussed in [24], there are more than 1500 papers on the WL algorithm as well as variants such as the $1/t$-WL algorithm [3]. These algorithms are not well understood; some variants are guaranteed to converge asymptotically [10], but bounds on convergence rate or accuracy seem to be lacking. For a representative example, see for example [23], which describes a Gibbs distribution model of protein folding, and uses the WL algorithm to determine relevant properties.
@@
We again emphasize that the continuous-setting algorithm can be used for integer distributions, and is often more efficient depending on the parameters. The problem ${P^{{\varepsilon}}_{\tt ratio-all}}$ was first considered in [16], under the name omnithermal approximation, along with an algorithm of cost $O(\frac{q^{2}}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$. A number of algorithms have been developed for pointwise estimation of $Q(\beta_{\max})$, which we denote by ${P^{{\varepsilon}}_{\tt ratio-pt}}$, with steadily improving sample complexities [4, 25, 14]; the best prior algorithm [14] had cost $O((q\log n)(\log q+\log\log n+\varepsilon^{-2}))$. No specialized algorithms were known for the integer setting.
We also show lower bounds for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ (and, a fortiori, for ${P^{{\varepsilon}}_{\tt ratio-all}}$) for black-box sampling algorithms, which match Theorem 3 up to $\operatorname*{polylog}(n,q)$ factors:
@@
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
  • $\mu$ is continuous and ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega\bigl{(}\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
  • $\mu$ is log-concave and ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega\bigl{(}\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}.$
@@
The definitions of the two problems ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, and especially the precise form of the error term for ${P^{\delta,\varepsilon}_{\tt count}}$, may seem somewhat arbitrary. To illustrate, we note a few reductions showing how the error bounds for these problems can be combined to give estimates of the entire family of Gibbs distribution:
@@
Given a solution for ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, we can produce estimates
$\hat{\mu}_{\alpha}(x)=\frac{\hat{\pi}(x)e^{\alpha x}}{\hat{Q}(\alpha\mid\mathcal{D})}$
for any $\alpha\in[\beta_{\min},\beta_{\max}],x\in\mathcal{F}$ such that $|\hat{\mu}_{\alpha}(x)-\mu_{\alpha}(x)|\leq 3\varepsilon(\mu_{\alpha}(x)+\delta)$.
@@
Indeed, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting and log-concave setting immediately give us the corresponding algorithms in Theorem 3, essentially for free. See Appendix A for proofs.
@@ 1.2 Algorithmic sampling-to-counting
As a concrete application of our algorithms, consider the following scenario: we have a combinatorial system with weighted items, and we have an algorithm to (approximately) sample from the corresponding Gibbs distribution. The runtime (e.g. mixing time of a Markov chain) may depend on $\beta$. As a few prominent examples:
@@
A closely related statistical problem is to count the objects of a given weight class. There have been a number of problem-specific methods to use the sampling algorithm for this task, e.g. [18, 9]. Our algorithms for ${P_{\tt count}}$ can be combined with any prior sampling algorithms to yield improved and unified algorithmic results, essentially for free. As some examples, we will show the following:
@@
Let $G=(V,E)$ be a graph and for each $i=0,\dots,|V|/2$ let $c_{i}$ be the number of matchings in $G$ with $i$ edges. Suppose $c_{|V|/2}>0$ and $c_{|V|/2-1}/c_{|V|/2}\leq f$ for a known parameter $f$. There is an FPRAS to simultaneously estimate all counts $c_{i}$ to relative error $\varepsilon$ in time $\tilde{O}(|E||V|^{3}f/\varepsilon^{2})$. In particular, if $G$ has minimum degree at least $|V|/2$, the time complexity is $\tilde{O}(|V|^{7}/\varepsilon^{2})$.
@@
Second, we may only have access to an approximate oracle $\tilde{\mu}_{\beta}$ instead of the exact oracle $\mu_{\beta}$. For example, given access to an MCMC sampler with stationary distribution $\mu_{\beta}$ and mixing time $\tau$, we can estimate $\mu_{\beta}$ to total variation distance $\rho$ by running for $\Theta(\tau\log(1/\rho))$ iterations. By a standard coupling argument, our results remain valid if exact oracles are replaced with sufficiently close approximate oracles (see e.g. [25, Remark 5.9]); see Appendix B for a formal statement.
@@ 1.3 Overview
We will develop two, quite distinct, types of algorithms: the first uses “cooling schedules” similar to [14], and the second is based on a new type of “covering schedule” for the integer setting. In Section 8, we show the lower bounds for the problems ${P_{\tt ratio-pt}}$ and ${P_{\tt count}}$.
Before the technical details, let us provide a high-level roadmap. For simplicity, we assume that tasks need to be solved with constant success probability.
@@
For the problem ${P^{{\varepsilon}}_{\tt ratio-all}}$, we extend a statistical method of [14] called the Paired Product Estimator (see Section 3). This constructs a cooling schedule $\beta_{0},\beta_{1},\ldots,\beta_{t-1},\beta_{t}$, where the log partition function is near-linear between successive values $\beta_{i}$. As shown in [14], this property allows us to successively estimate the ratios $Q(\beta_{i})/Q(\beta_{i-1})$ via an unbiased estimator with low variance. In addition, we show that this property allows us to fill in values $Q(\alpha)$ for $\alpha\in(\beta_{i},\beta_{i-1})$ via linear interpolation.
@@
In the log-concave setting, there is a more efficient algorithm for ${P^{\delta,\varepsilon}_{\tt count}}$. Here, for a fixed index $i$ of the covering schedule and for $j\in[k_{i},k_{i+1}]$ we have $\mu_{\beta_{i}}(j)\geq\min\{\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})\}=w_{i}$; thus, $\beta_{i}$ “covers” the interval $[k_{i},k_{i+1}]$. We can solve ${P^{\delta,\varepsilon}_{\tt count}}$ with $O(\frac{1}{\delta\varepsilon^{2}}+\sum_{i}\frac{1}{w_{i}\varepsilon^{2}})$ samples, by drawing $\Theta(\frac{1}{w_{i}\varepsilon^{2}})$ samples at $\beta_{i}$ and $\Theta(\frac{1}{\delta\varepsilon^{2}})$ samples at $\beta_{\min}$ and $\beta_{\max}$.
@@
First, we construct a “preschedule” where each interval can choose two different indices $\sigma^{-}_{i},\sigma^{+}_{i}$ instead of a single index $k_{i}$, with the indices interleaving as $\sigma_{i}^{-}\leq\sigma_{i+1}^{-}\leq\sigma_{i}^{+}\leq\sigma_{i+1}^{+}$. The algorithm repeatedly fill gaps: if some half-integer $\ell+1/2$ is not currently covered, we can select a value $\beta$ with $\mu_{\beta}([0,\ell])\approx\mu_{\beta}([\ell+1,n])$, along with corresponding values $\sigma^{-}\in[0,\ell],\sigma^{+}\in[\ell+1,n]$ with $\mu_{\beta}(\sigma^{+})\cdot(\sigma^{+}-\ell)\geq\Omega(\frac{1}{\log n})$ and $\mu_{\beta}(\sigma^{-})\cdot(\ell-\sigma^{-}+1)\geq\Omega(\frac{1}{\log n})$. The interval $[\sigma^{-},\sigma^{+}]$ then fills the gap and has weight $w\geq\Omega(\frac{1}{(\sigma^{+}-\sigma^{-})\log n})$.
@@ 2 Preliminaries
Define $z(\beta)=\log Z(\beta)$ and $z(\beta_{1},\beta_{2})=\log\frac{Z(\beta_{2})}{Z(\beta_{1})}=\log\frac{Q(\beta_{2})}{Q(\beta_{1})}$; note that $z(\beta_{\min},\beta_{\max})\leq q$ by definition. We write $z^{\prime}$ for the derivative of function $z$; observe (see [14]) that ${\mathbb{E}}_{X\sim\mu_{\beta}}[X]=z^{\prime}(\beta)$. As a consequence, $z^{\prime}$ is nonnegative everywhere and so the function $z$ is concave-up.
@@
We write $\mu_{\beta}(x,y),\mu_{\beta}[x,y)$ instead of $\mu_{\beta}((x,y)),\mu_{\beta}([x,y))$, etc. for readability. For any fixed value $\chi$, the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is a non-increasing function of $\chi$, since $\frac{\mu_{\beta}[\chi,n]}{1-\mu_{\beta}[\chi,n]}=\frac{\sum_{x\geq\chi}c_{x}e^{\beta x}}{\sum_{x<\chi}c_{x}e^{\beta x}}$; when $\beta$ increases by $\delta$, the numerator is multiplied by at least $e^{\delta\chi}$ and the denominator is multiplied by at most $e^{\delta\chi}$. At several places, we also make use of the following inequality:

### v11

@@ 1 Introduction
A central problem in computational statistics is to convert a procedure for sampling combinatorial objects into a procedure for counting those objects, and vice versa. In this work, we consider sampling algorithms for Gibbs distributions. Formally, given a real-valued function $H(\cdot)$ over a finite set $\Omega$, the Gibbs distribution is the family of distributions $\mu^{\Omega}_{\beta}$ over $\Omega$, parameterized by $\beta$ over an interval $[\beta_{\min},\beta_{\max}]$, of the form
@@
The normalizing constant $Z(\beta)=\sum_{\omega\in\Omega}e^{\beta H(\omega)}$ is called the partition function. These distributions appear in a number of sampling algorithms and are also common in physics, where the parameter $-\beta$ corresponds to the inverse temperature, and $H(\omega)$ is called the Hamiltonian of the system.
@@
We also consider an associated probability distribution we call the gross Gibbs distribution $\mu_{\beta}(x)$ over $\mathbb{R}$ given by
@@
When $c_{0}>0$, we allow the value $\beta=-\infty$; here $Z(-\infty)=c_{0}$ and $\mu_{-\infty}$ is simply the distribution which is equal to 00 with probability one. In particular, for $\beta_{\min}=-\infty$ we have $Q(\beta)=Z(\beta)/c_{0}$.
@@
To avoid degenerate cases, we always assume $n,q\geq 2$ and $\varepsilon,\gamma\in(0,\tfrac{1}{2})$ unless stated otherwise. If upper bounds $n$ and/or $q$ are not available directly, they can often be estimated by simple algorithms (up to constant factors), or can be guessed by exponential back-off strategies; for simplicity we do not discuss such issues here.
@@ 1.1 Our contributions
Our main contribution is a black-box algorithm to estimate counts. To explain this, we must address two technical issues. First, note that in the gross Gibbs distributions, the counts can only be recovered up to scaling. So some (arbitrary) normalization must be chosen; we use a convenient normalization $\pi(x)$ defined as:
@@
With these two provisos, we can define ${P^{\delta,\varepsilon}_{\tt count}}$ for parameters $\delta,\varepsilon\in(0,\tfrac{1}{2})$ to be the problem of finding a vector $\hat{\pi}\in[0,1]^{\mathcal{F}}$ with the following two properties:
@@
In particular, if $\Delta(x)\geq\delta$, then ${P^{\delta,\varepsilon}_{\tt count}}$ provides a $1\pm 3\varepsilon$ relative approximation to $\pi(x)$. We emphasize here that when the algorithm succeeds (which should have probability at least $1-\gamma$), this guarantee holds simultaneously for all values $x$. We develop three main algorithmic results:
@@
We also show lower bounds for black-box algorithms to solve ${P^{\delta,\varepsilon}_{\tt count}}$ by sampling the gross Gibbs distribution. We summarize these results as follows:
@@
Let $n>n_{0},q>q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0}$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
@@
Some specialized algorithms have been developed and analyzed for count estimation in combinatorial problems, see e.g. [18, 9]. Alternatively, the Wang-Landau (WL) algorithm [26] is a popular counting heuristic in physical applications; this uses a completely different methodology from our algorithm, based on a random walk on $\mathcal{F}$ with a running count estimate $\hat{c}$. As discussed in [24], there are more than 1500 papers on the WL algorithm and variants such as the $1/t$-WL algorithm [3]. Some variants are guaranteed to converge asymptotically [10], but bounds on convergence rate or accuracy seem to be lacking. For a representative example, see [23], which describes a Gibbs distribution to model protein folding, and uses the WL algorithm to determine relevant properties.
@@
We again emphasize that the continuous-setting algorithm can be used for integer distributions, and is often more efficient depending on the parameters. The problem ${P^{{\varepsilon}}_{\tt ratio-all}}$ was first considered in [16], under the name omnithermal approximation, along with an algorithm of cost $O(\frac{q^{2}}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$. A number of algorithms have been developed for pointwise estimation of $Q(\beta_{\max})$, which we denote by ${P^{{\varepsilon}}_{\tt ratio-point}}$, with steadily improving sample complexities [4, 25, 14]; the best prior algorithm [14] had cost $O((q\log n)(\log q+\log\log n+\varepsilon^{-2}))$. No specialized algorithms were known for the integer setting.
We also show lower bounds for ${P^{{\varepsilon}}_{\tt ratio-point}}$ (and, a fortiori, for ${P^{{\varepsilon}}_{\tt ratio-all}}$) for black-box sampling algorithms, which match Theorem 3 up to $\operatorname*{polylog}(n,q)$ factors:
@@
Let $n>n_{0},q>q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0}$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
  • $\mu$ is continuous and ${P^{{\varepsilon}}_{\tt ratio-point}}$ requires cost $\Omega\bigl{(}\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
  • $\mu$ is log-concave and ${P^{{\varepsilon}}_{\tt ratio-point}}$ requires cost $\Omega\bigl{(}\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}.$
@@
The definitions of the two problems ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, and especially the error term for ${P^{\delta,\varepsilon}_{\tt count}}$, may seem obscure. For motivation, we note a few ways to combine their error bounds for estimates of the entire family of Gibbs distributions. (See Appendix A for proofs.)
@@
Given solutions for ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, we can estimate $\mu_{\alpha}$ for any $\alpha\in[\beta_{\min},\beta_{\max}]$ and $x\in\mathcal{F}$ via:
$\hat{\mu}_{\alpha}(x)=\frac{\hat{\pi}(x)e^{\alpha x}}{\hat{Q}(\alpha\mid\mathcal{D})}\qquad\qquad\text{with $|\hat{\mu}_{\alpha}(x)-\mu_{\alpha}(x)|\leq 3\varepsilon(\mu_{\alpha}(x)+\delta)$}$
@@
Indeed, via Theorem 6, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting and log-concave setting immediately give us the corresponding algorithms in Theorem 3.
@@ 1.2 Algorithmic sampling-to-counting
As a concrete application of our algorithms, consider the following scenario: we have a combinatorial system with weighted items, along with an algorithm to (approximately) sample from the corresponding Gibbs distribution. The runtime (e.g. mixing time of a Markov chain) may depend on $\beta$. As a few prominent examples:
@@
Our algorithms for ${P_{\tt count}}$ can be combined with any prior sampling algorithms to yield improved and unified counting algorithms, essentially for free. As some examples, we will show the following:
@@
Let $G=(V,E)$ be a graph and let $v=|V|/2$. For each $i=0,\dots,v$ let $c_{i}$ be the number of matchings in $G$ with $i$ edges. Suppose $c_{v}>0$ and $c_{v-1}/c_{v}\leq f$ for a known parameter $f$. There is an FPRAS to simultaneously estimate all counts $c_{i}$ to relative error $\varepsilon$ in time $\tilde{O}(|E||V|^{3}f/\varepsilon^{2})$. In particular, if $G$ has minimum degree at least $|V|/2$, the time complexity is $\tilde{O}(|V|^{7}/\varepsilon^{2})$.
@@
Second, we may only have access to an approximate oracle $\tilde{\mu}_{\beta}$ instead of the exact oracle $\mu_{\beta}$. For example, given access to an MCMC sampler with stationary distribution $\mu_{\beta}$ and mixing time $\tau$, we can estimate $\mu_{\beta}$ to total variation distance $\rho$ by running for $\Theta(\tau\log\tfrac{1}{\rho})$ iterations. By a standard coupling argument, our results remain valid if exact oracles are replaced with sufficiently close approximate oracles (see e.g. [25, Remark 5.9]); see Appendix B for a formal statement.
@@ 1.3 Overview
We will develop two, quite distinct, types of algorithms: the first uses “cooling schedules” similar to [14], and the second is based on a new type of “covering schedule” for the integer setting. In Section 8, we show the lower bounds for the problems ${P_{\tt ratio-point}}$ and ${P_{\tt count}}$.
Let us provide a high-level roadmap here. For simplicity, we assume that tasks need to be solved with constant success probability.
@@
For the problem ${P^{{\varepsilon}}_{\tt ratio-all}}$, we extend a statistical method of [14] called the Paired Product Estimator (see Section 3). This constructs a cooling schedule $\beta_{0},\beta_{1},\ldots,\beta_{t-1},\beta_{t}$, where the log partition function is near-linear between successive values $\beta_{i}$. As shown in [14], this property gives an unbiased estimator with low variance for the ratios $Q(\beta_{i})/Q(\beta_{i-1})$. In addition, we show how to use this property to fill in values $Q(\alpha)$ for $\alpha\in(\beta_{i},\beta_{i-1})$ via linear interpolation.
@@
In the log-concave setting, there is a more efficient algorithm for ${P^{\delta,\varepsilon}_{\tt count}}$. Here, for any index $i$ of the covering schedule and for $j\in[k_{i},k_{i+1}]$ we have $\mu_{\beta_{i}}(j)\geq\min\{\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})\}=w_{i}$; thus, $\beta_{i}$ “covers” the interval $[k_{i},k_{i+1}]$. We can solve ${P^{\delta,\varepsilon}_{\tt count}}$ with $O(\frac{1}{\delta\varepsilon^{2}}+\sum_{i}\frac{1}{w_{i}\varepsilon^{2}})$ samples, by drawing $\Theta(\frac{1}{w_{i}\varepsilon^{2}})$ samples at $\beta_{i}$ and $\Theta(\frac{1}{\delta\varepsilon^{2}})$ samples at $\beta_{\min}$ and $\beta_{\max}$.
@@
First, we construct a “preschedule” where each interval can choose two different indices $\sigma^{-}_{i},\sigma^{+}_{i}$ instead of a single index $k_{i}$, with the indices interleaving as $\sigma_{i}^{-}\leq\sigma_{i+1}^{-}\leq\sigma_{i}^{+}\leq\sigma_{i+1}^{+}$. The algorithm repeatedly fill gaps: if some half-integer $\ell+\tfrac{1}{2}$ is not currently covered, we can select a value $\beta$ with $\mu_{\beta}([0,\ell])\approx\mu_{\beta}([\ell+1,n])$, along with corresponding values $\sigma^{-}\in[0,\ell],\sigma^{+}\in[\ell+1,n]$ with $\mu_{\beta}(\sigma^{+})\cdot(\sigma^{+}-\ell)\geq\Omega(\frac{1}{\log n})$ and $\mu_{\beta}(\sigma^{-})\cdot(\ell-\sigma^{-}+1)\geq\Omega(\frac{1}{\log n})$. The interval $[\sigma^{-},\sigma^{+}]$ then fills the gap and has weight $w\geq\Omega(\frac{1}{(\sigma^{+}-\sigma^{-})\log n})$.
@@ 2 Preliminaries
Define $z(\beta)=\log Z(\beta)$ and $z(\beta_{1},\beta_{2})=\log\frac{Z(\beta_{2})}{Z(\beta_{1})}=\log\frac{Q(\beta_{2})}{Q(\beta_{1})}$; note that $z(\beta_{\min},\beta_{\max})\leq q$ by definition. We write $z^{\prime}$ and $z^{\prime\prime}$ for the derivatives of function $z$; observe (see [14]) that ${\mathbb{E}}_{X\sim\mu_{\beta}}[X]=z^{\prime}(\beta)$. As a consequence, $z^{\prime},z^{\prime\prime}$ are nonnegative everywhere and the function $z$ is increasing concave-up.
@@
We write $\mu_{\beta}(x,y),\mu_{\beta}[x,y),\mu_{\beta}[x,y]$ instead of $\mu_{\beta}((x,y)),\mu_{\beta}([x,y)),\mu_{\beta}([x,y])$ for readability. For any fixed value $\chi$, the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is a non-increasing function of $\chi$, since $\frac{\mu_{\beta}[\chi,n]}{1-\mu_{\beta}[\chi,n]}=\frac{\sum_{x\geq\chi}c_{x}e^{\beta x}}{\sum_{x<\chi}c_{x}e^{\beta x}}$; when $\beta$ increases by $\delta$, the numerator is multiplied by at least $e^{\delta\chi}$ and the denominator is multiplied by at most $e^{\delta\chi}$. At several places, we also make use of the following inequality: